OpenStudy (j2lie):

9a^2+6a-15 My answer is 3a(a+5)-3(3a+5) I don't know what I did wrong.

OpenStudy (anonymous):

missing a \(3\)

OpenStudy (j2lie):

oh oops.

OpenStudy (j2lie):

Thank you.

OpenStudy (anonymous):

actually more is wrong

OpenStudy (usukidoll):



OpenStudy (j2lie):

?

OpenStudy (j2lie):

is there anything else I did wrong? @agent0smith

OpenStudy (shamil98):

3(a-1)(3a+5)

OpenStudy (usukidoll):

o crap sorry didn't distribute the negative oy yes you can factor it.

OpenStudy (j2lie):

Thank you. I did that.

OpenStudy (shamil98):

9a^2+6a-15 = (3a-3)(3a+5) = 3(a-1)(3a+5)

OpenStudy (usukidoll):